the length of a rectangular room is 9 feet longer than twice the width. if the room's perimeter is 186ft what are the room's dimensons?
Answer by haileytucki(390)

2(w)+2(w+9)=186
Multiply 2 by each term inside the parentheses.
2w+2(w+9)=186
Multiply 2 by each term inside the parentheses.
2w+2w+18=186
Since 2w and 2w are like terms, add 2w to 2w to get 4w.
4w+18=186
Since 18 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 18 from both sides.
4w=-18+186
Add 186 to -18 to get 168.
4w=168
Divide each term in the equation by 4.
(4w)/(4)=(168)/(4)
Simplify the left-hand side of the equation by canceling the common factors.
w=(168)/(4)
Simplify the right-hand side of the equation by simplifying each term.
width=42 length=50